# BoundMatch: Boundary detection applied to semi-supervised segmentation for urban-driving scenes

### v3

@@ 4.4 More Real-World SS-SS Setting
Inspired by the evaluation protocol in [49], we assess SS-SS in a realistic scenario: using all 2,975 Cityscapes training images as labeled data and the additional 19,997 “extra” images as truly unlabeled data. In this setting, the fully supervised baseline is already strong, making further gains challenging yet highly relevant for practical deployment. As shown in Tab. 11, SAMTH+BoundMatch achieves consistent improvements over both SAMTH and UniMatch, particularly in boundary-specific metrics (BIoU and BF1), demonstrating its effectiveness at leveraging large unlabeled pools in real-world applications.
@@ 4.5 Application to lightweight models
Table 12: Lightweight segmentation architectures applied to the same setting as Tab. 11. Here “BoundMatch” refers to SAMTH+BoundMatch which is trained with the additional unlabeled Extra data. We use input size of $768\times 768$ and $800\times 800$ for MobileNet-V2 and AFFormer-Tiny, respectively.
@@
DLV3+ MV2 Baseline $73.4$ $53.9$ $56.5$ $89.7$G $5.8$M $36.0$
BoundMatch $78.2$ $59.0$ $64.3$ $93.7$G $6.0$M $31.8$
AFFormer-Tiny Baseline $76.7$ $57.5$ $65.2$ $7.3$G $2.1$M $89.1$
Mobile-Seed $78.0$ $58.3$ $65.1$ $10.0$G $2.4$M $70.2$
BoundMatch $80.1$ $61.1$ $68.2$ $9.2$G $2.2$M $78.0$
@@
We further evaluate BoundMatch on lightweight segmentation architectures suitable for resource-constrained devices. Specifically, we apply our method to a MobileNet-V2 backbone [64] with DeepLabV3+ and to AFFormer-Tiny, an efficient Transformer model [22]. For the latter, we compare against Mobile-Seed [12], a strong baseline for boundary-aware lightweight segmentation using AFFormer-Tiny.
Results in Tab. 12 demonstrate that applying BoundMatch yields substantial gains in both overall accuracy and boundary quality metrics across both lightweight backbones. These accuracy improvements are achieved with only a minor increase in computational overhead (FLOPs and inference time), indicating a favorable trade-off between performance and efficiency compared to the original lightweight models.
Notably, when applied to AFFormer-Tiny, BoundMatch not only surpasses the segmentation quality of the specialized Mobile-Seed method but also achieves higher throughput (FPS) with slightly lower FLOPs. This highlights the efficiency benefits of our MTL architecture combined with SS-SS training on additional unlabeled data, even compared to mobile segmentation methods.
These findings confirm the BoundMatch framework’s flexibility and its effectiveness in enhancing performance even for models designed for real-time, low-compute deployment scenarios.
5 Conclusion and Future Work
In this paper, we introduced BoundMatch, a novel framework that addresses the critical challenge of imprecise boundaries in semi-supervised semantic segmentation. BoundMatch integrates semantic boundary detection into consistency regularization through three key components: (1) BCRM, which enforces consistency on both segmentation masks and semantic boundaries derived from hierarchical features; (2) BSF, which injects learned boundary cues into the segmentation decoder; and (3) SGF, which refines boundary predictions using spatial gradients for higher-quality pseudo-labels. Designed as a flexible framework, BoundMatch can enhance various SS-SS baselines, as demonstrated through successful integration with SAMTH, UniMatch, and PrevMatch.
Our extensive experiments across five diverse datasets demonstrate that BoundMatch consistently improves performance when applied to different baselines, achieving state-of-the-art results with DINOv2 foundation models. Systematic evaluation using boundary-specific metrics (BIoU and BF1) reveals boundary quality improvements that traditional mIoU fails to capture. Additionally, we introduced SAMTH, a strong baseline featuring HBN that addresses training instabilities in teacher-student frameworks with EMA updates.
Future work will explore extending BoundMatch’s boundary-aware learning principles to 3D perception tasks and investigating tighter integration with vision-language models. Our results demonstrate that explicit boundary modeling, when properly integrated through dedicated consistency regularization and fusion mechanisms, significantly enhances segmentation quality; particularly crucial for applications requiring precise object delineation such as autonomous driving and robotic manipulation.
@@ Appendix A Harmonious Batch Normalization (HBN)
  2. 2.
    Generate hard pseudo-labels: $\hat{p}^{T}\leftarrow\operatorname*{argmax}p^{T,w}_{c}$ (apply confidence $\tau$ if needed).
  3. 3.
    Student forward pass (train mode, twin-view): $p^{S,L},p^{S,w},p^{S,s}\leftarrow f_{\theta^{S}}(x,u^{w},u^{s})$; also update student BN buffers with $\rho$.
  4. 4.
    Loss computation & student update: compute $\mathcal{L}^{L}$ on $(p^{S,L},y)$ and $\mathcal{L}^{U}$ on $(p^{S,s},\hat{p}^{T})$; update $\theta^{S}$.
  5. 5.
@@
In teacher-student consistency training the teacher weights $\theta^{T}$ are updated as an exponential moving average (EMA) of the student weights $\theta^{S}$. Copying the student’s batch normalization (BN) running statistics, common in prior work, makes those statistics incompatible with $\theta^{T}$, because the teacher’s parameter trajectory and input distribution differ from the student’s. HBN fixes this by
@@ B.1 Boundary Detection Head Design
In Fig. 9, we illustrate the architecture of the boundary detection head used in our BoundMatch framework. We follow CASENet [26] and use a multi-scale boundary detection head that consists of four “Side Layers”. Each “Side Layer” consists of a $1\times 1$ convolution up sampling to $\nicefrac{{1}}{{2}}$ of the input image size, followed by a $3\times 3$ convolution similar to the one introduced in SBCB [11]. For more details on the boundary detection head, please refer to the original paper [26].
@@ B.2 DINOv2 with BoundMatch
In Fig. 10, we illustrate the architecture of DPT with boundary detection head used in our BoundMatch framework. Following the design of DPT, we use the reassembled feature maps used for the Fusion modules as the hierarchical features for the boundary detection head. For more details on DPT, please refer to the original paper [51].
@@
(a) Per-class IoU results.
(b) Per-class BIoU results.
(c) Per-class BF1 results.
Figure 11: Per-class performance between SAMTH and SAMTH+BoundMatch on Cityscapes dataset using $\nicefrac{{1}}{{16}}$ split and ResNet-50 backbone. Green values represent improvements while red values indicate degradation.
Figure 14: Qualitative results on the Pascal VOC 2012 dataset on the difficult Classic 92 image split. We compare SAMTH and SAMTH+BoundMatch. Best viewed zoomed in and in color.
Figure 15: Qualitative results on the Cityscapes dataset on $\nicefrac{{1}}{{16}}$ split. We compare SAMTH and SAMTH+BoundMatch using DPT with DINOv2-S encoder. Best viewed zoomed in and in color.
Appendix C Boundary Evaluation Metrics
@@
BIoU evaluates the overlap between predicted ($\hat{y}$) and ground-truth segmentation masks $y$ restricted only to pixels lying within a narrow-band around object boundaries. Formally, BIoU is computed as
@@
where $BD_{k}(y)=y\oplus\mathrm{MinPool}_{k}(y)$ denotes the binary boundary mask obtained by applying a $k\times k$ min-pooling (stride 1) to the ground-truth segmentation map $y$ and taking the pixel-wise XOR ($\oplus$) with the original mask, and $\wedge$ and $\vee$ denote pixel-wise logical AND and OR, respectively [65]. Because it ignores interior pixels, BIoU directly measures boundary alignment and penalizes both over- and under-segmentation along object edges. It is therefore particularly sensitive to boundary precision but agnostic to interior region accuracy. We use a 5 pixel radius ($k=11$) for the boundary mask in our experiments.
@@
[…]
where precision and recall are computed by matching predicted and ground-truth boundary pixels within a small tolerance radius [66]. Formally, precision and recall can be defined as:
@@
BF1 thus provides insight into the trade-off between missing fine boundary details (low recall) and producing spurious edges (low precision), while allowing for minor spatial misalignments. We use a 5 pixel tolerance radius ($k=11$) for our experiments.
@@
BIoU emphasizes exact boundary overlap, making it well suited for tasks requiring crisp, well-aligned boundaries. BF1, by explicitly balancing precision and recall under a spatial tolerance, is more robust to slight shifts but sensitive to fragmented or noisy boundary predictions. Reporting both metrics alongside mIoU therefore provides a holistic view of a model’s segmentation performance, particularly its ability to delineate object boundaries accurately.
Appendix D Per-Class Performance
Here we present the per-class performance comparison between SAMTH and BoundMatch for Cityscapes $\nicefrac{{1}}{{16}}$ split in Figs. 11(a), 11(b) and 11(c) for IoU, BIoU, and BF1 metrics respectively. For IoU, we observe that BoundMatch improves the performance on most of the classes, especially on vehicles and pedestrians. However, BoundMatch drastically degrades the performance on the fence category, which is understandable because the fence category is annotated with a filled silhouette that covers both the rails and all the gaps between them. Perhaps, the model focuses on segmenting the objects behind the fence, which leads to a low IoU score. An example of this happening is shown in Fig. 12.
The per-class BIoU and BF1 results in Figs. 11(b) and 11(c) show that BoundMatch improves the boundary quality of most of the classes, especially on vehicles and pedestrians again. There are no categories with significant degradation, which indicates that BoundMatch is able to improve the boundary quality of most of the categories.
Appendix E More Qualitative Results
E.1 BDD100K

### v4

@@ 4.4 More Real-World SS-SS Setting
Inspired by the evaluation protocol in [49], we assess SS‑SS in a realistic scenario: using all 2,975 Cityscapes training images as labeled data and the additional 19,997 “extra” images as truly unlabeled data. In this setting, the fully supervised baseline is already strong, making further gains challenging yet highly relevant for practical deployment. As shown in Tab.˜12, SAMTH+BoundMatch achieves consistent improvements over both SAMTH and UniMatch, particularly in boundary‑specific metrics (BIoU and BF1), demonstrating its effectiveness at leveraging large unlabeled pools in real‑world applications.
@@ 4.5 Application to lightweight models
Table 13: Lightweight segmentation architectures applied to the same setting as Tab.˜12. Here “BoundMatch” refers to SAMTH+BoundMatch which is trained with the additional unlabeled Extra data. We use input size of $768\times 768$ and $800\times 800$ for MobileNet-V2 and AFFormer-Tiny, respectively.
@@
DLV3+ MV2 Baseline $73.4$ $53.9$ $56.5$ 89.789.7G 5.85.8M $36.0$
BoundMatch $78.2$ $59.0$ $64.3$ 93.793.7G 6.06.0M $31.8$
AFFormer-Tiny Baseline $76.7$ $57.5$ $65.2$ 7.37.3G 2.12.1M $89.1$
Mobile-Seed $78.0$ $58.3$ $65.1$ 10.010.0G 2.42.4M $70.2$
BoundMatch $80.1$ $61.1$ $68.2$ 9.29.2G 2.22.2M $78.0$
@@
We evaluate BoundMatch on lightweight architectures suitable for resource-constrained deployment, specifically DeepLabV3+ with MobileNet-V2 [66] and AFFormer-Tiny [22]. For AFFormer-Tiny, we compare against Mobile-Seed [12], a state-of-the-art boundary-aware lightweight segmentation method. More information about the experimental setup is provided in Appendix B.3.
As shown in Tab.˜13, BoundMatch achieves accuracy improvements of 3.4–4.8% mIoU with moderate impact on runtime performance, maintaining above 30 FPS for real-time applications. For MobileNet-V2, we achieve 4.8% IoU and 7.8% boundary F1 gains with only 4.5% additional FLOPs, maintaining 31.8 FPS suitable for real-time applications. More notably, AFFormer-Tiny with BoundMatch reaches 80.1% IoU and 61.1% BIoU while sustaining 78 FPS, well above the 30 FPS threshold required for real-time video processing. With only 2.2M parameters and 9.2 GFLOPs, this configuration is particularly suited for mobile deployment where memory and power constraints are important considerations.
Compared to the specialized Mobile-Seed method, BoundMatch achieves superior accuracy with 8% fewer FLOPs and 11% higher throughput, demonstrating that our multi-task learning framework with semi-supervised training provides a more efficient solution than dedicated boundary architectures. The modest computational overhead (approximately 26% additional FLOPs for AFFormer-Tiny) is justified by boundary quality improvements important for applications such as autonomous navigation, medical imaging on portable devices, and augmented reality systems.
These results confirm that BoundMatch scales to lightweight architectures without compromising deployment feasibility, making it practical for enhancing segmentation quality in resource-constrained environments where both accuracy and efficiency are important.
4.6 Limitations and Future Work
While BoundMatch demonstrates consistent improvements across benchmarks, several limitations merit discussion and guide future research:
Challenging structures: Classes with fine, partially transparent, or heavily occluded structures (e.g. fences, see-through facades) can degrade performance, likely due to annotation policies (filled silhouettes) and the model’s tendency to favor occluded content behind meshes (Fig.˜10 and LABEL:fig:app_failure_cases). Future work should explore attention-based modules for long-range context modeling and investigate alternative supervision strategies that better model transparency and occlusion.
[…]
Domain generalization: BoundMatch shows strong performance across most evaluated domains including remote-sensing (Tab.˜20), but does not show gains on medical imaging (Tab.˜19). The boundary detection components developed in this work have not been optimized for medical imaging settings, which may require different design considerations. Future work could investigate how to adapt boundary-aware learning to diverse application domains with their specific requirements and BoundMatch may benefit from dataset-specific optimizations.
Annotation quality dependencies: BoundMatch’s performance is fundamentally limited by the quality of the underlying segmentation annotations. Errors or inconsistencies in segmentation masks propagate through both the training process and boundary generation via distance transforms. This dual dependency is particularly problematic in datasets with known annotation issues (e.g. Pascal VOC’s ignore regions and Cityscapes’ inconsistent annotations LABEL:fig:app_failure_cases) where both segmentation and derived boundaries inherit the same biases. Future work should explore self-supervised boundary discovery or leverage foundation models’ priors to reduce dependence on precise annotations. Additionally, developing robust training strategies that explicitly account for annotation noise in both tasks could improve real-world deployment.
Modern architecture integration: As vision transformers and foundation models become prevalent, adapting BoundMatch’s principles presents opportunities. Incorporating boundary-aware consistency into vision-language models could enhance their spatial understanding for tasks requiring precise localization. The multi-task consistency principle could extend to other complementary views (depth, surface normals) in semi-supervised settings, particularly as foundation models provide increasingly rich priors for such tasks [52].
5 Conclusion
This paper presented BoundMatch, a framework that integrates semantic boundary detection into semi-supervised segmentation through consistency regularization. Rather than deriving boundaries from segmentation outputs, we learn them from hierarchical features through independent task heads, with BSF and SGF modules enabling bidirectional information flow between tasks.
Our experiments across six datasets show consistent improvements of 0.4–2.4% mIoU alongside gains in boundary-specific metrics (BIoU and BF1). The modular design allows flexible deployment—using all components or selecting based on efficiency requirements. Additionally, our Harmonious Batch Normalization (HBN) addresses training instabilities in teacher-student frameworks, improving baseline methods by up to 16.4% mIoU.
BoundMatch integrates with existing methods (UniMatch, PrevMatch), scales to modern architectures (DINOv2 transformers), and maintains benefits on lightweight models for real-time deployment. While improvements are incremental, their consistency validates boundary-aware learning as a useful principle for semi-supervised segmentation, particularly for applications requiring precise object delineation such as autonomous driving.
Looking forward, this multi-task consistency approach could extend to other geometric cues from foundation models, while incorporating boundary awareness into vision-language models could enhance their spatial precision. Understanding the theoretical basis for boundary-segmentation complementarity remains an important open question.
@@ Appendix A Harmonious Batch Normalization (HBN)
  2. 12.
    Generate hard pseudo-labels: $\hat{p}^{T}\leftarrow\operatorname*{argmax}p^{T,w}_{c}$ (apply
confidence $\tau$ if needed).
  • 23.
    Student forward pass (train mode, twin-view): $p^{S,L},p^{S,w},p^{S,s}\leftarrow f_{\theta^{S}}(x,u^{w},u^{s})$; also update student BN buffers
  • with $\rho$.
  • 34.
    Loss computation & student update: compute $\mathcal{L}^{L}$ on $(p^{S,L},y)$ and
  • $\mathcal{L}^{U}$ on $(p^{S,s},\hat{p}^{T})$; update $\theta^{S}$.
  • 5.
@@
    In teacher-student consistency training the teacher weights $\theta^{T}$ are updated as an exponential moving average (EMA) of the student weights $\theta^{S}$. Copying the student’s batch normalization (BN) running statistics, common in prior work, makes those statistics incompatible with $\theta^{T}$, because the teacher’s parameter trajectory and input distribution differ from those of the student. HBN fixes this by
@@ B.1 Boundary Detection Head Design
    In Fig.˜11, we illustrate the architecture of the boundary detection head used in our BoundMatch framework. We follow CASENet [26] and use a multi-scale boundary detection head that consists of four “Side Layers”. Each “Side Layer” consists of a $1\times 1$ convolution up sampling to $\nicefrac{{1}}{{2}}$ of the input image size, followed by a $3\times 3$ convolution similar to the one introduced in SBCB [11]. For more details on the boundary detection head, please refer to the original paper [26].
@@ B.2 DINOv2 with BoundMatch
    In Fig.˜12, we illustrate the architecture of DPT with boundary detection head used in our BoundMatch framework. Following the design of DPT, we use the reassembled feature maps used for the Fusion modules as the hierarchical features for the boundary detection head. For more details on DPT, please refer to the original paper [51].
@@
    B.3 Lightweight Models with BoundMatch
    For DeepLabV3+ with MobileNet-V2 [66], we attach the boundary detection head to the output of the backbone in the same manner as DeepLabV3+ with ResNet backbones. The hyperparameters used for training are identical to the setup introduced in Sec.˜4.1.
    For integrating BoundMatch to AFFormer-Tiny [22], we follow Mobile-Seed [12] and use four features from the backbone (Blocks 1, 3, 5, and 6). The input to the decoder assumes input channel size of 216 and the feature channel size before the classification layer is 96, which is identical to Mobile-Seed and AFFormer. For feature fusion with BSF, we use one convolutional layer which takes in concatenated features from the backbone and boundary detection head to obtain the final feature. The training hyperparamters are identical to the one used in AFFormer [22].
    Appendix C Boundary Region Evaluation
    C.1 Boundary Evaluation Metrics
@@
    BIoU evaluates the overlap between predicted ($\hat{y}$) and ground‑truth segmentation masks $y$ restricted only to pixels lying within a narrow-band around object boundaries. Formally, BIoU is computed as
@@
    where $BD_{k}(y)=y\oplus\mathrm{MinPool}_{k}(y)$ denotes the binary boundary mask obtained by applying a $k\times k$ min‑pooling (stride 1) to the ground‑truth segmentation map $y$ and taking the pixel‑wise XOR ($\oplus$) with the original mask, and $\wedge$ and $\vee$ denote pixel‑wise logical AND and OR, respectively [67]. Because it ignores interior pixels, BIoU directly measures boundary alignment and penalizes both over‑ and under‑segmentation along object edges. It is therefore particularly sensitive to boundary precision but agnostic to interior region accuracy. We use a 5 pixel radius ($k=11$) for the boundary mask in our experiments.
@@
    $\mathrm{BF1}\;=\;\frac{1}{C}\sum_{c=1}^{C}\frac{2\,\mathrm{Prec}_{c}\,\mathrm{Rec}_{c}}{\mathrm{Prec}_{c}+\mathrm{Rec}_{c}},$ (20)
    where precision and recall are computed by matching predicted and ground‑truth boundary pixels within a small tolerance radius [58]. Formally, precision and recall can be defined as:
@@
    BF1 thus provides insight into the trade‑off between missing fine boundary details (low recall) and producing spurious edges (low precision), while allowing for minor spatial misalignments. We use a 5 pixel tolerance radius ($k=11$) for our experiments.
@@
    BIoU emphasizes exact boundary overlap, making it well suited for tasks requiring crisp, well‑aligned boundaries. BF1, by explicitly balancing precision and recall under a spatial tolerance, is more robust to slight shifts but sensitive to fragmented or noisy boundary predictions. Reporting both metrics alongside mIoU therefore provides a holistic view of a model’s segmentation performance, particularly its ability to delineate object boundaries accurately.
    C.2 BIoU and BF1 Evaluation of Benchmark Results
    Table 14: Component analysis of the proposed modules on Pascal VOC 2012 val set using ResNet-50 backbone. We report the mean IoU, Boundary IoU (BIoU), and Boundary F1 (BF1) scores. The best results are highlighted in bold.
    $\nicefrac{{1}}{{16}}$ $\nicefrac{{1}}{{8}}$
    BCRM BSF SGF IoU BIoU BF1 IoU BIoU BF1
    $73.5$ $67.9$ $59.0$ $75.8$ $68.7$ $59.2$
    $75.6$ $68.3$ $59.2$ $76.8$ $69.1$ $59.3$
    $75.9$ $68.1$ $59.3$ $76.5$ $69.3$ $60.0$
    $75.8$ $68.0$ $59.2$ $76.6$ $68.8$ $59.8$
    76.3 68.9 59.8 77.2 69.7 60.3
    Table 15: Boundary Metrics for Cityscapes Benchmark.
    ResNet-50 ResNet-101
    $\nicefrac{{1}}{{16}}$ $\nicefrac{{1}}{{8}}$ $\nicefrac{{1}}{{16}}$ $\nicefrac{{1}}{{8}}$
    BIoU BF1 BIoU BF1 BIoU BF1 BIoU BF1
    SAMTH $54.8$ $59.3$ $56.8$ $61.8$ $55.8$ $61.4$ $57.9$ $63.1$
    + BoundMatch $56.6$ $63.2$ $57.8$ $64.7$ $58.3$ $65.2$ $59.4$ $66.7$
    UniMatch $55.3$ $60.7$ $56.6$ $63.5$ $56.9$ $63.3$ $57.8$ $65.1$
    + BoundMatch $56.0$ $62.6$ $57.5$ $64.2$ $57.6$ $64.8$ $58.8$ $66.4$
    PrevMatch $55.3$ $60.7$ $57.0$ $63.3$ $57.5$ $63.7$ $58.8$ $65.6$
    + BoundMatch $56.8$ $63.4$ $58.2$ $64.4$ $58.0$ $65.4$ $58.9$ $66.4$
    Table 16: Boundary Metrics for Pascal VOC Classic Split.
    ResNet-50 ResNet-101
    92 183 92 183
    BIoU BF1 BIoU BF1 BIoU BF1 BIoU BF1
    SAMTH $65.2$ $55.0$ $65.6$ $53.4$ $70.7$ $59.5$ $70.5$ $59.6$
    + BoundMatch $67.2$ $56.9$ $68.3$ $57.6$ $70.9$ $61.3$ $72.7$ $62.5$
    UniMatch $64.9$ $55.2$ $65.2$ $56.1$ $69.3$ $57.4$ $69.9$ $60.1$
    + BoundMatch $68.2$ $57.7$ $68.6$ $57.5$ $70.2$ $59.0$ $72.1$ $61.7$
    PrevMatch $67.1$ $57.1$ $67.3$ $57.2$ $69.8$ $60.7$ $71.8$ $61.8$
    + BoundMatch $69.0$ $58.0$ $68.7$ $57.4$ $71.0$ $61.2$ $72.4$ $61.8$
    Table 17: Boundary Metrics for Pascal VOC Blender Split.
    $\nicefrac{{1}}{{16}}$ $\nicefrac{{1}}{{8}}$
    BIoU BF1 BIoU BF1
    SAMTH $68.5$ $59.8$ $69.2$ $59.1$
    + BoundMatch $68.8$ $60.1$ $70.0$ $59.9$
    UniMatch $66.6$ $57.3$ $68.0$ $57.2$
    + BoundMatch $69.4$ $59.5$ $70.4$ $60.4$
    PrevMatch $68.4$ $57.7$ $69.3$ $58.2$
    + BoundMatch $68.7$ $59.1$ $69.4$ $59.3$
    Table 18: Boundary Metrics for BDD100K, SYNTHIA, and ADE20K.
    BDD100K SYNTHIA ADE20K
    $\nicefrac{{1}}{{64}}$ $\nicefrac{{1}}{{32}}$ $\nicefrac{{1}}{{64}}$ $\nicefrac{{1}}{{32}}$ $\nicefrac{{1}}{{128}}$ $\nicefrac{{1}}{{64}}$
    BIoU BF1 BIoU BF1 BIoU BF1 BIoU BF1 BIoU BF1 BIoU BF1
    Supervised $31.8$ $34.2$ $35.9$ $37.4$ $55.3$ $68.7$ $60.3$ $74.8$ $5.8$ $6.9$ $7.6$ $8.3$
    UniMatch $37.9$ $42.5$ $41.1$ $45.5$ $60.8$ $74.5$ $63.6$ $78.1$ $9.2$ $10.2$ $12.7$ $12.8$
    Ours $39.8$ $46.1$ $41.3$ $47.3$ $62.2$ $76.7$ $64.8$ $80.1$ $12.0$ $13.0$ $14.9$ $16.2$
    In Tabs.˜15, 16, 17 and 18, we show the boundary evaluation metrics (BIoU and BF1) for each of the datasets used in the main paper. While comprehensive boundary evaluation of all prior methods would be ideal, most existing SS-SS works do not provide open-source implementations or pretrained models, making it infeasible to fairly compute these specialized metrics retrospectively. We therefore present boundary metrics for the representative methods we could reliably reproduce (SAMTH, UniMatch, PrevMatch) alongside their BoundMatch-enhanced versions. Despite this limitation, the results clearly demonstrate that BoundMatch consistently improves boundary quality across different datasets and baseline models, validating our approach’s effectiveness in enhancing boundary delineation.
    Appendix D Additional Ablation Studies
    D.1 Component Analysis of Pascal VOC
    In Tab.˜14, we show the component analysis of our BoundMatch framework on Pascal VOC Classic 92 with ResNet-50 backbone. We can see that each component of BoundMatch contributes to the overall performance. The most contribution comes from the Boundary-aware Consistency Regularization Module (BCRM), which improves over 1.0% compared to the SAMTH baseline. Incremental, but gradual improvements can be seen when introducing BSF and SGF, and results in over 1.4–2.8% improvements.
    D.2 Pseudo-Label Accuracy
    Figure 13: (a) and (b) presents pseudo-label accuracies (mIoU) of the validation set against training iterations for Cityscapes and Pascal VOC Classic 92, respectively. We compare SAMTH, UniMatch, and PrevMatch with and without BoundMatch.
    In Fig.˜13, we plot the pseudo-label accuracy (IoU) on the validation set against training iterations for Cityscapes and Pascal VOC Classic 92. The training curves reveal BoundMatch’s noise-robustness properties. On Cityscapes LABEL:fig:a, all methods (with and without BoundMatch) show monotonically increasing pseudo-label accuracy without late-stage degradation, indicating that consistency regularization prevents overfitting to early noisy pseudo-labels. On Pascal VOC LABEL:fig:b, where boundary annotations are often noisier, BoundMatch notably stabilizes the fluctuating pseudo-label accuracy observed in baseline methods, particularly for UniMatch and PrevMatch. SAMTH+BoundMatch shows improvements over SAMTH in terms of accuracy, but overfitting can be seen after 40K iterations, indicating the need for stronger regularizations which are equipped in UniMatch and PrevMatch.
    This stabilization can be attributed to the multi-task learning framework: when segmentation pseudo-labels become noisy, the boundary task, learned from hierarchical features and refined through SGF, provides a strong regularization signal to deter the model in overfitting. This aligns with the noise-robust learning principles identified in [37], where combining multiple complementary views and regularization mechanisms proves more effective than single approaches.
    Appendix E Additional Qualitative Results
    Figure 14: More qualitative results which compare SAMTH and SAMTH+BoundMatch on the Cityscapes dataset with $\nicefrac{{1}}{{16}}$ protocol: (a) improved cases and (b) failure cases. Best viewed zoomed in and in color. Note that we cropped and scaled each samples differently across each samples and boundary predictions could seem thicker due to being zoomed in.
    Figure 16: Qualitative results on the Pascal VOC 2012 dataset on the difficult Classic 92 image split. We compare SAMTH and SAMTH+BoundMatch. The white region in the ground truth masks are “ignore” regions. Best viewed zoomed in and in color.
    Figure 17: Qualitative results on the Cityscapes dataset on $\nicefrac{{1}}{{16}}$ split. We compare SAMTH and SAMTH+BoundMatch using DPT with DINOv2-S pretrained ViT encoder. Best viewed zoomed in and in color.
    Table 19: Results on the ACDC Dataset. The results are reported in Dice Similarity Coefficient.
    Cases
    11 33 77
    Supervised $28.5$ $41.5$ $62.5$
    CPS [43] $61.0$ $81.5$
    UniMatch [9] $85.4$ $88.9$ $89.9$
    UniMatch + BoundMatch $85.0$ $89.0$ $89.6$
    Table 20: Results on the LoveDA dataset.
    $\nicefrac{{1}}{{32}}$ $\nicefrac{{1}}{{16}}$
    IoU BIoU BF1 IoU BIoU BF1
    Supervised $45.2$ $27.0$ $24.8$ $45.7$ $29.1$ $25.3$
    UniMatch $48.9$ $28.7$ $28.3$ $48.8$ $29.4$ $29.7$
    SAMTH $49.8$ $29.7$ $29.8$ $50.1$ $30.7$ $29.6$
    SAMTH + BoundMatch $51.1$ $30.3$ $30.8$ $52.1$ $31.3$ $31.1$
    E.1 Cityscapes
    In LABEL:fig:app_successful_cases, we present qualitative comparisons between SAMTH and SAMTH+BoundMatch on the Cityscapes dataset. For each example, we display the segmentation masks, prediction errors, and semantic boundary predictions (for SAMTH+BoundMatch only). BoundMatch reduces segmentation errors near object boundaries, as demonstrated in examples (1) through (5). Notably, BoundMatch better preserves object continuity: in (3) and (4), it accurately segments complete "terrain", "rider", and "truck" regions rather than producing fragmented results. Despite the inherent challenge of segmenting thin structures, BoundMatch successfully captures "poles" in examples (1) and (5), where the baseline typically struggles. The strong alignment between predicted boundaries and actual object edges indicates that the segmentation head leverages boundary information to generate more accurate masks.
    In LABEL:fig:app_failure_cases, we show notable failure cases for SAMTH+BoundMatch on the Cityscapes dataset. We believe thin objects like "poles" show limited improvement for two main reasons: first, small objects create imbalanced supervision signals, and second, annotation inconsistencies lead to false positives. For instance, images (6) and (7) show the same scene from different samples, but poles are annotated only in (7), not in (6). This inconsistency causes false positives, particularly for SAMTH+BoundMatch, which aims to precisely segment boundaries. BoundMatch also struggles with challenging scenarios like reflections, as seen in (8). Additionally, transparent or semi-transparent structures pose difficulties: in (9) and (10), see-through "fences" and "buildings" challenge BoundMatch as it attempts to accurately segment partially occluded objects behind them.
    E.2 BDD100K